y(x) =(sinx) linx


1
Expert's answer
2021-05-27T06:04:28-0400
"y=(\sin x)^{\ln x}"

Then


"\ln y=\ln{(\sin x)^{\ln x}}"

"\ln y=\ln x\cdot\ln{(\sin x)}"

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}(\ln y)=\dfrac{d}{dx}(\ln x\cdot\ln{(\sin x)})"

Use the Chain Rule


"\dfrac{1}{y}\cdot\dfrac{dy}{dx}=\dfrac{\ln (\sin x)}{x}+\dfrac{\ln x}{\sin x}\cdot \cos x"

"\dfrac{dy}{dx}=y(\dfrac{\ln (\sin x)}{x}+\dfrac{\ln x}{\tan x})"

"\dfrac{dy}{dx}=(\sin x)^{\ln x}\bigg(\dfrac{\ln (\sin x)}{x}+\dfrac{\ln x}{\tan x}\bigg)"